HETAV PATEL

Grade 11,

How to solve this 0.55x200 upon 11 and the answer is 10 how it is solved

How to solve this
 
0.55x200 upon 11
and the answer is 10
how it is solved
 

Grade:8

4 Answers

Carmel Emmanuel
8 years ago
first multiply 55/100 x 200 which gives 110 now divide by 11 final ans is 10. 
Mondo Camry
8 years ago
just multipy the nmbers then didvide it
Nitesh
6 years ago
After removing decimal from 0.55 we get 55/100 Now multiply with 200 we get 55×200/100=110.Now divide this with 11 that is 110/11=10 and so 10 is the answer.
Siddharth Singh
6 years ago
55/100 *200 =110...........110/11 =10................Thanksgiving from Siddharth Singh at askiitians .com